OpenStudy (anonymous):

(x^2 + 3)^2 - 6(x^2 + 3) = 7

OpenStudy (anonymous):

let (x^2 + 3) = t and solve for t then solve (x^2 + 3 = t for all values of t

OpenStudy (anonymous):

A sub looks good, maybe a difference of squares, too.

OpenStudy (anonymous):

Oh, that's stupid, it's not squared, lol..

OpenStudy (anonymous):

x^4-16=0 (X^2+4)(x^2-4)=0 (x^2+4)(x-2)(x+2) x=2i,-2i,2,-2

OpenStudy (anonymous):

dont forget to resub in t and you will get my answer